# Massively Parallel Minimum Spanning Tree in General Metric Spaces

### v1

@@ 2.1 Technical Overview of Theorem 1
    In parallel for all $t=1/\alpha,1,\alpha,\alpha^{2},\ldots,W$, construct $\widehat{\mathcal{P}}_{2t},\widehat{\mathcal{P}}_{4t},\widehat{\mathcal{P}}_{8t},\ldots,\widehat{\mathcal{P}}_{\alpha t/2}$: The problem is solved in a divide-and-conquer manner. In the first iteration, $\widehat{\mathcal{P}}_{\alpha t}$ is given and the goal is to compute $\widehat{\mathcal{P}}_{\sqrt{\alpha}t}$. One can start with merging each cluster in the auxiliary partition $\mathcal{P}^{\prime}_{\sqrt{\alpha}t}$ as a super-node and compute connected components on the graph $G_{\leq\sqrt{\alpha}t}$ after removing the edges that cross $\widehat{\mathcal{P}}_{\alpha t}$. $\widehat{\mathcal{P}}_{\sqrt{\alpha}t}$ is composed of the connected components found by the above process. Then the problem requires recursively computing: $(1)$ $\widehat{\mathcal{P}}_{2t},\widehat{\mathcal{P}}_{4t},\ldots,\widehat{\mathcal{P}}_{\sqrt{\alpha}t/2}$ and $(2)$ $\widehat{\mathcal{P}}_{2\sqrt{\alpha}t},\widehat{\mathcal{P}}_{4\sqrt{\alpha}t},\ldots,\widehat{\mathcal{P}}_{\alpha t/2}$.
Two major challenges to implementing the above algorithm in the MPC model are (1) computing the auxiliary hierarchy $\{\mathcal{P}^{\prime}_{i}\}$, (2) computing the connected components. For challenge (1), [jayaram2024massively] solved it by applying randomly shifted grids over Euclidean points. For challenge (2), they observed that each connected component subroutine starts with super-nodes (incomplete connected components) obtained by merging clusters of $\mathcal{P}^{\prime}_{t}$ for some $t$, and all points within the same connected component are within distance $t\cdot\textnormal{poly}(\log n)$. In this case, they showed that one only needs to reduce the number of incomplete connected components by a $\textnormal{poly}(\log n)$ factor, and connecting the remaining incomplete connected components in an arbitrary way that does not increase the final MST cost by too much. By applying standard leader compression based connected components algorithms [andoni2018parallel, behnezhad2019near, liu2020connected, karloff2010model, reif1984optimal], each connected component subroutine only takes $O(\log\log n)$ rounds. Since the second phase has $O(\log\log\log n)$ levels of recursions, the overall number of rounds is $O(\log\log(n)\cdot\log\log\log(n))$.
@@
[…]
Another challenge is the second phase of the algorithm in [jayaram2024massively] which is both inefficient (incurs an additional $\log\log\log n$ factor on the running-time) and results in a large constant approximation instead of our desired $1+\varepsilon$ approximation. In the following, we first explain how to handle the first phase for input general metric points, and then describe our novel second-phase algorithm.
Let us consider the complete metric graph $G$ over $V$, i.e., each edge $(u,v)$ has weight $w(u,v)$. To start, we utilize the low-diameter decomposition [MillerPX13] of $G$. It is computed as follows: (1) Add an additional source node, and add an edge from the source node to each node in $G$ with some random weight. (2) Compute a single source shortest path tree from the added source node and make each subtree of the source node a cluster of the decomposition. Since every shortest path in $G$ only uses one hop, the depth of the shortest path tree is always at most $2$, and thus it can be easily computed by $O(1)$ rounds of the Bellman-Ford algorithm. As a result, we get a sequence of partitions $\{\mathcal{P}^{0}_{\alpha^{k}}\}_{k}$ with similar guarantees to those of the shifted grids. That is, on level $t$, each cluster has diameter at most $t$, and the crossing probability of an edge $(u,v)$ is $w(u,v)\cdot\textnormal{poly}(\log n)/t$.
We need to make sure the partitions form a hierarchy. To achieve this, we construct $\mathcal{P}^{\textnormal{\tiny MPX}}=\{\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}\}_{k}$ as follows. For each level $t=\alpha^{k}$, the partition $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ is obtained by intersecting all the low-diameter decompositions $\mathcal{P}^{0}_{t^{\prime}}$ for $t^{\prime}\geq t$, i.e., $u$ and $v$ are in the same cluster of $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ iff $u$ and $v$ are in the same cluster of $\mathcal{P}^{0}_{t^{\prime}}$ for $t^{\prime}\geq t$. Notice that this does not increase the diameter of each cluster. In addition, the crossing probability of an edge $(u,v)$ is still bounded by the geometric sum $\sum_{t^{\prime}\geq t}w(u,v)\cdot\textnormal{poly}(\log n)/t^{\prime}$ which is at most $\frac{1}{1-\frac{1}{\alpha}}\leq 2$ times the original probability. We need to further process $\mathcal{P}^{\textnormal{\tiny MPX}}$ in order to lower the crossing probabilities by a factor of $\alpha$, and obtain the final hierarchy $\widehat{\mathcal{P}}=\{\widehat{\mathcal{P}}_{\alpha^{k}}\}_{k}$.
@@
The final hierarchy $\widehat{\mathcal{P}}=\{\widehat{\mathcal{P}}_{\alpha^{k}}\}_{k}$ is obtained by extending the partition of $\mathcal{P}^{\textnormal{\tiny MPX}}$. The partition $\widehat{\mathcal{P}}_{t}$ is meant to approximate the set of connected components with respect to the edges of weight at most $t$ that do not cross $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$, call these edges $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$ (see the figure on the right). For every level $t=\alpha^{k}$, $\widehat{\mathcal{P}}_{t}$ is obtained by executing $\Theta\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds of leader compression on $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$, using the edges of $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$. After the leader compression, the “incomplete” components within each cluster of $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$ are joined together, where we say a component is incomplete when it still has outgoing edges that would have been used if we had not stopped the leader compression early.
For the second phase, we develop a completely new algorithm. The original phase-2 algorithm aimed to generate intermediary partitions $\widehat{\mathcal{P}}_{2\cdot t},\widehat{\mathcal{P}}_{4\cdot t},\widehat{\mathcal{P}}_{8\cdot t},\ldots,\widehat{\mathcal{P}}_{\alpha\cdot t/2}$, between $\widehat{\mathcal{P}}_{t}$ and $\widehat{\mathcal{P}}_{\alpha t}$. The final approximate MST was obtained by applying a spanning tree algorithm for each level. There are two major disadvantages: (1) Generating all intermediate partitions $\widehat{\mathcal{P}}_{t^{\prime}}$ introduces an additional $O(\log\log\log n)$ rounds overhead. (2) Computing an arbitrary spanning tree for each level introduces a factor of $2$ in the approximation ratio. To address these two issues, we forgo generating the intermediary partitions. Instead, we use a novel variation Borůvka’s algorithm to build the approximate MST.
For each level $t=\alpha^{k}$, we regard each cluster of $\widehat{\mathcal{P}}_{t/\alpha}$ as a super-node and try to find a minimum spanning forest over the edges of weight at most $\alpha t$ that do not cross $\widehat{\mathcal{P}}_{t}$, call these edges $G_{\leq\alpha t}[\widehat{\mathcal{P}}_{t}]$. That is, the forest connects all the super-nodes inside one cluster of $\widehat{\mathcal{P}}_{t}$. This way, we avoid computing the intermediate levels, and it saves us the $O(\log\log\log n)$ factor in the number of MPC rounds. We propose the following variant of Borůvka’s algorithm:
@@
Had we run this algorithm for $\Theta(\log n)$ rounds, the exact minimum forest with respect to $G_{\leq\alpha t}[\widehat{\mathcal{P}}_{t}]$ would have been found. As it turns out, using many fewer rounds, i.e. $T=\Theta\left(\frac{\alpha}{\varepsilon}\right)$ computes most of the minimum spanning forest, and hence suffices for a $(1+\varepsilon)$-approximation. Consider level $t=\alpha^{k}$. Let $C$ be an arbitrary cluster of $\widehat{\mathcal{P}}_{t}$ that is composed of clusters $C_{1},C_{2},\ldots,C_{r}$ of $\widehat{\mathcal{P}}_{t/\alpha}$. Let us apply the above algorithm to super-nodes $C_{1},\ldots C_{r}$ over the edges $G_{\leq\alpha t}[\widehat{\mathcal{P}}_{t}]$. At any point, unless $C_{1},\ldots,C_{r}$ are merged into one super-node, during the next round, each super-node joins another one with a constant probability. Therefore, after $T$ rounds, we expect the number of unmerged super-nodes in $C$ to be (roughly) $r/2^{\Theta(T)}$. That is, a $\left(\textnormal{poly}\left(\frac{\varepsilon}{\alpha}\right)\right)$-fraction of the edges of the minimum spanning forest are already discovered. We connect the remaining super-nodes in $C$ using arbitrary edges of weight at most $\alpha t$. This is fine, as we have assured there are not many super-nodes left inside $C$.
Observe that the approximation ratio is controlled by $\alpha$. a larger $\alpha$ improves the approximation ratio by introducing lower crossing probabilities for the edges, whereas a smaller $\alpha$ means more parallelization. Specifically, (approximately) computing the minimum spanning forest for $\widehat{\mathcal{P}}_{t/\alpha}$ requires $\Omega\left(\log\frac{\alpha}{\varepsilon}\right)$ rounds. To achieve a $(1+\varepsilon)$-approximation ratio, we need to set $\alpha=\textnormal{poly}\left(\frac{\alpha}{\varepsilon}\right)$, which results in an $O\left(\log\frac{1}{\varepsilon}+\log\log n\right)$-round algorithm.
2.2 Technical Overview of Theorem 2
One of our main contributions is a $\Omega\left(\log\frac{1}{\varepsilon}\right)$-round lower bound for computing a $(1+\varepsilon)$-approximate of the MST in $(1,2)$-metrics (Theorem 2). In fact, our lower bound even applies to the $(1,2)$-metrics where the weight-$1$ edges form disjoint cycles. We consider only such metrics for this discussion. An approximate MST excludes at least one weight-$1$ edge from each cycle, and (on average) not many more. Specifically, if the weight-$1$ edges form two cycles, one of length $\frac{1}{10\varepsilon}$, and one of length $n-\frac{1}{10\varepsilon}$, then a $(1+\varepsilon)$-approximate MST excludes at least one edge from each cycle, and excludes at most a total of $2\varepsilon n$ edges.
@@ 2.2 Technical Overview of Theorem 2
This intuition can be formalized to obtain a $\Omega\left(\frac{1}{\varepsilon}\right)$ lower bound in the distributed LOCAL model where the nodes are allowed to communicate over the weight-$1$ edges of the metric (see Section A.1). Then, using the framework of [GhaffariKU19] and [CzumajDP21a] it can be lifted to obtain a (conditional) $\Omega\left(\log\frac{1}{\varepsilon}\right)$ lower bound for component stable MPC algorithms, where the components are with respect to weight-$1$ edges (see Section A.2). That is, the output of the MPC algorithm on a connected component of the weight-$1$ edges is required to be independent of the rest of the graph. While many MPC algorithms are component stable, many natural ones are not, as discussed in [CzumajDP21a]. We show that the assumption on component stability can be dropped. Specifically, in Theorem 2, we give a lower bound assuming solely the 1vs2-Cycle conjecture.
@@
The $\frac{1}{10\varepsilon}$-neighborhood of a vertex is detected as follows. Assume for now that the input graph is a cycle of length $n$. Given two vertices $u_{1}$ and $u_{2}$, the goal is to assert whether they are at most $\frac{1}{10\varepsilon}$ hops away in the cycle. We start by picking one adjacent edge for each of $u_{1}$ and $u_{2}$, and doing a two-switch (see Figure 1). If $u_{1}$ and $u_{2}$ are close, and the appropriate edges are picked, this results in the cycle being broken down into two cycles: one of length at most $\frac{1}{10\varepsilon}$ containing $u_{1}$ and $u_{2}$, and another of length at least $n-\frac{1}{10\varepsilon}$ containing. We consider the $(1,2)$-metric where the weight-$1$ edges are these two cycles, and study the probabilities of each edge being excluded when $\mathcal{A}^{\prime}$ is run on the metric. It can be shown that the edges in the smaller cycle are excluded with probability at least $10\varepsilon$, and the edges in the larger cycle with probability at most $2\varepsilon$. Therefore, they can be identified, and we can assert that $u_{1}$ and $u_{2}$ are close.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$(a)$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$(b)
Figure 1: A two-switch on edges $(u,v)$ and $(u^{\prime},v^{\prime})$, replaces them with $(u,u^{\prime})$ and $(v,v^{\prime})$. Here, the two-switch breaks the cycle into two cycles. See also Definition 5.4 and Claim 5.5.
Using the $\frac{1}{10\varepsilon}$-neighborhoods, we aim to decrease the length of the cycle by a factor of $\varepsilon$. After doing this $O(\log_{1/\varepsilon}n)$ times, we are able to verify whether the original graph was a cycle of length $n$. A potential approach is to select certain vertices as leaders and contract their $\frac{1}{10\varepsilon}$ neighborhoods. It is not immediately clear this is possible in $O(1)$ rounds, since we need at most $O(\varepsilon n)$ leaders such that each vertex joins a leader, and the vertices that join the same leader form a connected subgraph (i.e. a subpath) of the cycle. Instead of contracting, we break down the cycle into $\frac{1}{\varepsilon}$ smaller cycles by removing all the edges, and connecting each vertex to the two vertices that were exactly $\frac{1}{\varepsilon}$ hops away in the cycle (Figure 2). This can be done repeatedly. However, the subroutine for computing the $\frac{1}{10\varepsilon}$-neighborhoods becomes more complicated.

### v2

@@ 2.1 Technical Overview of Theorem˜1
    In parallel for all $t=1/\alpha,1,\alpha,\alpha^{2},\ldots,W$, construct $\widehat{\mathcal{P}}_{2t},\widehat{\mathcal{P}}_{4t},\widehat{\mathcal{P}}_{8t},\ldots,\widehat{\mathcal{P}}_{\alpha t/2}$: The problem is solved in a divide-and-conquer manner. In the first iteration, $\widehat{\mathcal{P}}_{t}$ and $\widehat{\mathcal{P}}_{\alpha t}$ are given and the goal is to compute $\widehat{\mathcal{P}}_{\sqrt{\alpha}t}$. One can start with contracting each vertex set in the auxiliary partition $\mathcal{P}^{\prime}_{\sqrt{\alpha}t}$ as a super-node and compute connected components on the graph $G_{\leq\sqrt{\alpha}t}$ after removing the edges that cross $\widehat{\mathcal{P}}_{\alpha t}$. $\widehat{\mathcal{P}}_{\sqrt{\alpha}t}$ is composed of the connected components found by the above process. Then the problem requires recursively computing: $(1)$ $\widehat{\mathcal{P}}_{2t},\widehat{\mathcal{P}}_{4t},\ldots,\widehat{\mathcal{P}}_{\sqrt{\alpha}t/2}$ and $(2)$ $\widehat{\mathcal{P}}_{2\sqrt{\alpha}t},\widehat{\mathcal{P}}_{4\sqrt{\alpha}t},\ldots,\widehat{\mathcal{P}}_{\alpha t/2}$.
Two major challenges arise when implementing the above algorithm in the MPC model: (1) computing the auxiliary hierarchy $\{\mathcal{P}^{\prime}_{i}\}$, (2) computing the connected components. For the former, [JMNZ24] solved it by applying randomly shifted grids over Euclidean points. For the latter, they observed that each connected component subroutine starts with super-nodes (incomplete connected components) obtained by merging vertex sets of $\mathcal{P}^{\prime}_{t}$ for some $t$, and all points within the same connected component are within distance $t\cdot\textnormal{poly}(\log n)$. In this case, they showed that one only needs to reduce the number of incomplete connected components by a factor of $\textnormal{poly}(\log n)$, and connecting the remaining incomplete connected components in an arbitrary way that does not increase the final MST cost by too much. By applying standard leader compression based connected components algorithms [ASSWZ18, BDELM19, LTZ20, KSV10, Rei84], each connected component subroutine only takes $O(\log\log n)$ rounds. Since the second phase has $O(\log\log\log n)$ levels of recursions, the overall number of rounds is $O(\log\log(n)\cdot\log\log\log(n))$.
@@
A major challenge for general metrics is to compute the auxiliary hierarchy of partitions. While [JMNZ24] use randomly shifted hypergrids in the Euclidean space, there is no natural data-oblivious hierarchical partitioning of general metrics. To get around this, we carefully implement a low-diameter decomposition of [MPX13] (henceforth MPX) in our setting.
Another issue with the algorithm in [JMNZ24] is the second phase. It is inefficient in terms of both the round-complexity and the approximation ratio. That is, it incurs an additional $\log\log\log n$ factor on the running-time, and it results in a large constant approximation instead of our desired $1+\varepsilon$ approximation ratio. In the following, we first explain how to handle the first phase for when the input is a general metric, and then describe our novel second-phase algorithm.
Let us consider the complete metric graph $G$ over $V$, i.e., each edge $(u,v)$ has weight $w(u,v)$. To start, we utilize the low-diameter decomposition [MPX13] of $G$. For a given parameter $t$, it can be computed as follows: (1) Add an additional source node, and add an edge from the source to each vertex in $G$. The weight of each edge is drawn from the exponential distribution with mean $t/\log n$. (2) Compute a single-source shortest-path tree from the source node. The vertex sets in the decomposition are the subtrees directly below the root. Since the weights in $G$ come from a metric, the depth of this tree is always at most $2$, and thus it can be easily computed in $O(1)$ rounds of the MPC model (e.g. using the Bellman-Ford algorithm).
We compute an MPX decomposition for each $t=\alpha^{k}$, where $\alpha=\Theta\left(\frac{\log^{2}n}{\varepsilon}\right)$, up to $t=W$. As a result, we get a sequence of partitions $\{\mathcal{P}^{0}_{\alpha^{k}}\}_{k}$ with similar guarantees to those of the shifted grids. That is, on level $t$, each vertex set has diameter at most $t$, and the crossing probability of an edge $(u,v)$ is $w(u,v)\cdot\textnormal{poly}(\log n)/t$.
We need to make sure the partitions form a hierarchy. To achieve this, we construct $\mathcal{P}^{\textnormal{\tiny MPX}}=\{\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}\}_{k}$ as follows. For each level $t=\alpha^{k}$, the partition $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ is obtained by intersecting all the decompositions $\mathcal{P}^{0}_{t^{\prime}}$ for $t^{\prime}\geq t$, i.e., $u$ and $v$ are in the same set of $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ iff $u$ and $v$ are in the same set of $\mathcal{P}^{0}_{t^{\prime}}$ for all $t^{\prime}\geq t$. Notice that this does not increase the diameter of each set. In addition, the crossing probability of an edge $(u,v)$ is still bounded by the geometric sum $\sum_{t^{\prime}\geq t}w(u,v)\cdot\textnormal{poly}(\log n)/t^{\prime}$ which is at most $\frac{1}{1-\frac{1}{\alpha}}\leq 2$ times the original probability. We need to further process $\mathcal{P}^{\textnormal{\tiny MPX}}$ in order to lower the crossing probabilities by a factor of $\alpha$, and obtain the final hierarchy $\widehat{\mathcal{P}}=\{\widehat{\mathcal{P}}_{\alpha^{k}}\}_{k}$.
@@
The final hierarchy $\widehat{\mathcal{P}}=\{\widehat{\mathcal{P}}_{\alpha^{k}}\}_{k}$ is obtained by extending the partitions of $\mathcal{P}^{\textnormal{\tiny MPX}}$. The partition $\widehat{\mathcal{P}}_{t}$ is meant to approximate the set of connected components with respect to the edges of weight at most $t$ that do not cross $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$, call these edges $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$ (see the figure on the right). For every level $t=\alpha^{k}$, $\widehat{\mathcal{P}}_{t}$ is obtained by executing $\Theta\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds of leader compression on $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$, using the edges of $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$. After the leader compression, the “incomplete” components within each vertex set of $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$ are joined together, where we say a component is incomplete when it still has outgoing edges that would have been used if we had not stopped the leader compression early.
For the second phase, we develop a completely new algorithm. The original phase-2 algorithm aimed to generate intermediary partitions $\widehat{\mathcal{P}}_{2\cdot t},\widehat{\mathcal{P}}_{4\cdot t},\widehat{\mathcal{P}}_{8\cdot t},\ldots,\widehat{\mathcal{P}}_{\alpha\cdot t/2}$, between $\widehat{\mathcal{P}}_{t}$ and $\widehat{\mathcal{P}}_{\alpha t}$. The final approximate MST was obtained by computing an arbitrary spanning tree for each level. There are two major disadvantages to this: (1) Generating all intermediate partitions $\widehat{\mathcal{P}}_{t^{\prime}}$ introduces an additional $O(\log\log\log n)$-round overhead. (2) Computing an arbitrary spanning tree for each level introduces a factor of $2$ in the approximation ratio. To address these two issues, we forgo generating the intermediary partitions. Instead, we use a novel variation Borůvka’s algorithm to build the approximate MST.
For each level $t=\alpha^{k}$, we regard each vertex set of $\widehat{\mathcal{P}}_{t/\alpha}$ as a super-node and try to find a minimum spanning forest over the edges of weight at most $\alpha t$ that do not cross the next level partition $\widehat{\mathcal{P}}_{t}$, call these edges $G_{\leq\alpha t}[\widehat{\mathcal{P}}_{t}]$. That is, the forest connects all the super-nodes inside a set of $\widehat{\mathcal{P}}_{t}$. This way, we avoid computing the intermediate levels, and it saves us the $O(\log\log\log n)$ factor in the number of rounds. We propose the following variant of Borůvka’s algorithm:
@@
Had we run this algorithm for $\Theta(\log n)$ rounds, the exact minimum forest with respect to $G_{\leq\alpha t}[\widehat{\mathcal{P}}_{t}]$ would have been found. As it turns out, using many fewer rounds, i.e. $T=\Theta\left(\log\frac{\alpha}{\varepsilon}\right)$ computes most of the minimum spanning forest, and hence suffices for a $(1+\varepsilon)$-approximation. Consider level $t=\alpha^{k}$. Let $C$ be an arbitrary vertex set of $\widehat{\mathcal{P}}_{t}$ that is composed of sets $C_{1},C_{2},\ldots,C_{r}$ of $\widehat{\mathcal{P}}_{t/\alpha}$. Let us apply the above algorithm to super-nodes $C_{1},\ldots C_{r}$ over the edges $G_{\leq\alpha t}[\widehat{\mathcal{P}}_{t}]$. At any point, unless $C_{1},\ldots,C_{r}$ are merged into one super-node, during the next round, each super-node joins another one with a constant probability. Therefore, after $T$ rounds, we expect the number of unmerged super-nodes in $C$ to be (roughly) $r/2^{\Theta(T)}$. That is, only a $\left(\textnormal{poly}\left(\frac{\varepsilon}{\alpha}\right)\right)$-fraction of the minimum spanning forest edges remain undiscovered after $T$ rounds. We connect the remaining super-nodes in $C$ using arbitrary edges of weight at most $\alpha t$. This is fine, as we have assured there are not many super-nodes left inside $C$.
Observe that the approximation ratio is controlled by $\alpha$. A larger $\alpha$ improves the approximation ratio by introducing lower crossing probabilities for the edges, whereas a smaller $\alpha$ means more parallelization. Specifically, (approximately) computing the minimum spanning forest for $\widehat{\mathcal{P}}_{t/\alpha}$ requires $\Omega\left(\log\frac{\alpha}{\varepsilon}\right)$ rounds. To achieve a $(1+\varepsilon)$-approximation ratio, we need to set $\alpha=\Theta\left(\frac{\log^{2}n}{\varepsilon}\right)$, which results in an $O\left(\log\frac{1}{\varepsilon}+\log\log n\right)$-round algorithm.
2.2 Technical Overview of Theorem˜2
One of our main contributions is a $\Omega\left(\log\frac{1}{\varepsilon}\right)$-round lower bound for computing a $(1+\varepsilon)$-approximate of the MST in $(1,2)$-metrics (Theorem˜2). In fact, our lower bound even applies to the $(1,2)$-metrics where the weight-$1$ edges form disjoint cycles. We consider only such metrics for this discussion. An approximate MST excludes at least one weight-$1$ edge from each cycle, and (on average) not many more. Specifically, if the weight-$1$ edges form two cycles, one of length $\frac{1}{10\varepsilon}$, and one of length $n-\frac{1}{10\varepsilon}$, then a $(1+\varepsilon)$-approximate MST excludes at least one edge from each cycle, and excludes at most a total of $2\varepsilon n$ edges.
@@ 2.2 Technical Overview of Theorem˜2
This intuition can be formalized to obtain a $\Omega\left(\frac{1}{\varepsilon}\right)$ lower bound in the distributed LOCAL model where the nodes are allowed to communicate over the weight-$1$ edges of the metric (see Section˜A.1). Then, using the framework of [GKU19] and [CDP21] it can be lifted to obtain a (conditional) $\Omega\left(\log\frac{1}{\varepsilon}\right)$ lower bound for component-stable MPC algorithms, where the components are with respect to weight-$1$ edges (see Section˜A.2). That is, the output of the MPC algorithm on a connected component of the weight-$1$ edges is required to be independent of the rest of the graph. While many MPC algorithms are component-stable, many natural ones are not, as discussed in [CDP21]. We show that the assumption on component stability can be dropped. Specifically, in Theorem˜2, we give a lower bound assuming solely the 1vs2-Cycle conjecture.
@@
The $\frac{1}{10\varepsilon}$-neighborhood of a vertex is detected as follows. Assume for now that the input graph is a cycle of length $n$. Given two vertices $u_{1}$ and $u_{2}$, the goal is to assert whether they are at most $\frac{1}{10\varepsilon}$ hops away in the cycle. We start by picking one adjacent edge for each of $u_{1}$ and $u_{2}$, and doing a two-switch (see Figure˜1). If $u_{1}$ and $u_{2}$ are close, and the appropriate edges are picked, this results in the cycle being broken down into two cycles: one of length at most $\frac{1}{10\varepsilon}$ containing $u_{1}$ and $u_{2}$, and another of length at least $n-\frac{1}{10\varepsilon}$ containing. We consider the $(1,2)$-metric where the weight-$1$ edges are these two cycles, and study the probabilities of each edge being excluded when $\mathcal{A}^{\prime}$ is run on the metric. It can be shown that the edges in the smaller cycle are excluded with probability at least $10\varepsilon$, and the edges in the larger cycle with probability at most $2\varepsilon$. Therefore, they can be identified, and we can assert that $u_{1}$ and $u_{2}$ are close.
[…]
Figure 1: A two-switch on edges $(u,v)$ and $(u^{\prime},v^{\prime})$, replaces them with $(u,u^{\prime})$ and $(v,v^{\prime})$. Here, the two-switch breaks the cycle into two cycles. See also Definition˜5.4 and Claim˜$5.$
Using the $\frac{1}{10\varepsilon}$-neighborhoods, we aim to decrease the length of the cycle by a factor of $\varepsilon$. After doing this $O(\log_{1/\varepsilon}n)$ times, we are able to verify whether the original graph was a cycle of length $n$. A potential approach is to select certain vertices as leaders and contract their $\frac{1}{10\varepsilon}$ neighborhoods. It is not immediately clear this is possible in $O(1)$ rounds, since we need at most $O(\varepsilon n)$ leaders such that each vertex joins a leader, and the vertices that join the same leader form a connected subgraph (i.e. a subpath) of the cycle. Instead of contracting, we break down the cycle into $\frac{1}{\varepsilon}$ smaller cycles by removing all the edges, and connecting each vertex to the two vertices that were exactly $\frac{1}{\varepsilon}$ hops away in the cycle (Figure˜2). This can be done repeatedly. However, the subroutine for computing the $\frac{1}{10\varepsilon}$-neighborhoods becomes more complicated.